# Graph Neural Networks Gone Hogwild

### v1

@@ F.1 Architecture details
For GSDGNN (where Equation 7 describes the optimization objective for obtaining embeddings), we parameterize $g_{\theta}$ as an MLP with layer sizes $(16,16,16,2)$. We use the symmetric renormalized Laplacian matrix $\tilde{{\bm{L}}}={\bm{I}}-\tilde{{\bm{A}}}={\bm{I}}-({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}({\bm{A}}+{\bm{I}})({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}$ for the Laplacian regularization term, and set $\gamma=1.0,\beta=5.0$. (Note that gradient-based optimization of this objective function has a direct correspondence to the embedding update function of APPNP Gasteiger et al. (2019).)
@@ F.2 Training details
For binary classification experiments, we use binary cross entropy loss for training. For regression experiments, we use mean squared error. We use the Adam optimizer with weight decay, where we set the optimizer parameters as $\alpha=0.001,\beta_{1}=0.9,\beta_{2}=0.999$. We set the learning rate to $0.002$, and use exponential decay with rate $0.98$ ever $200$ epochs. We train for a maximum of 5000 epochs. These experiments were performed on a single NVIDIA RTX 2080 Ti.
@@ G.1 Dataset Details
The benchmark datasets we report performance for are MUTAG and PROTEINS, where the prediction task is graph classification, and PPI, where the prediction task is node classification.
@@
MUTAG is a dataset consisting of 188 graphs, each of which corresponds to a nitroaromatic compound (Srinivasan et al., 1996). The goal is to predict the mutagenicity of each compound on Salmonella typhimurium. Nodes in the graphs correspond to atoms (and are associated with a one-hot encoded feature in $\mathbb{R}^{7}$ corresponding to the atom type), and edges correspond to bonds. The average number of nodes in a graph is 17.93, and the average number of edges is 19.79.
@@
The PROTEINS dataset Borgwardt et al. (2005) consists of 1113 graphs, each of which corresponds to a protein. The task is predicting whether or not the protein is an enzyme. Nodes in the graph correspond to amino acids in the protein (and are associated with node features in $\mathbb{R}^{3}$ representing amino acid properties). Edges connect amino acids that are less than some threshold distance from one another in the protein. The average number of nodes in a graph is 39.06, and the average number of edges is 72.82.
@@
The PPI dataset (Hamilton, 2020) consists of 24 graphs, each of which corresponds to a protein-protein interaction network found in different areas of the body. Each node in the graph corresponds to a protein, with edges connecting proteins that interact with one another. Nodes are associated with features in $\mathbb{R}^{50}$, representing some properties of the protein. Each protein has $121$ binary prediction targets, each of which corresponds to some ontological property that the protein may or may not have. We use a 20/2/2 train/valid/test split consistent with Hamilton et al. (2017).
@@ G.2 Results
For PROTEINS and MUTAG, We perform 10-fold cross validation and report average classification accuracy and standard deviations in Table 3. We include performance reported by Gu et al. (2020) (marked by an asterisk), as well as performance of our own implementation of IGNN with a single layer and GCN with two layers. For PPI, we use a 20/2/2 train/valid/test split consistent with Hamilton et al. (2017).
Table 4 shows average micro-f1 scores for energy GNNs compared to other GNN architectures. We include performance reported by Gu et al. (2020) (marked by an asterisk). Where layer specifications are not included (for asterisked values), we were unable to determine them from Gu et al. (2020).
@@
GCN* (5 layer) 85.6 $\pm$ 5.8 76.0 $\pm$ 3.2
IGNN* (3 layer) 89.3 $\pm$ 6.7 77.7 $\pm$ 3.4
Energy GNN 87.6 $\pm$ 3.4 72.5 $\pm$ 0.3
@@
Table 3: Graph classification accuracy (%). Results are averaged (and standard deviations are computed) using 10 fold cross validation with 5 random parameter seeds. Asterisked values are from Gu et al. (2020).
@@
MLP* 46.2
GCN* 59.2
GraphSAGE* 78.6
GAT* (3 layer) 97.3
IGNN* (5 layer) 97.6
Energy GNN 76.2
@@
Table 4: Mean micro-F1 score on PPI dataset (%). Asterisked values are from Gu et al. (2020).

### v2

@@ F.1 Architecture details
For GSDGNN (where Equation 8 describes the optimization objective for obtaining embeddings), we parameterize $g_{\theta}$ as an MLP with layer sizes $(16,16,16,2)$. We use the symmetric renormalized Laplacian matrix $\tilde{{\bm{L}}}={\bm{I}}-\tilde{{\bm{A}}}={\bm{I}}-({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}({\bm{A}}+{\bm{I}})({\bm{D}}+{\bm{I}})^{-\frac{1}{2}}$ for the Laplacian regularization term, and set $\gamma=1.0,\beta=5.0$. (Note that gradient-based optimization of this objective function has a direct correspondence to the embedding update function of APPNP Gasteiger et al. [2019].)
@@ F.2 Training details
For binary classification experiments, we use binary cross entropy loss for training. For regression experiments, we use mean squared error. We use the Adam optimizer with weight decay, where we set the optimizer parameters as $\alpha=0.001,\beta_{1}=0.9,\beta_{2}=0.999$. We set the learning rate to $0.002$, and use exponential decay with rate $0.98$ ever $200$ epochs. In the forward pass for IGNN, we iterate on the node update equation until convergence of node embeddings, with a convergence tolerance of $10^{-5}$. The maximum number of iterations is set to 500. In the forward pass for the optimization-based GNNs, we use L-BFGS to minimize $E_{\theta}$ w.r.t node embeddings, with a convergence tolerance of $10^{-5}$. The maximum number of iterations is set to 50. We train for a maximum of 5000 epochs. These experiments were performed on a single NVIDIA RTX 2080 Ti.
@@ G.1 Dataset Details
The benchmark datasets we report performance for are MUTAG, PROTEINS, Peptides-func, and Peptides-struct, where the prediction task is graph classification/regression, and PPI, where the prediction task is node classification.
@@
MUTAG is a dataset consisting of 188 graphs, each of which corresponds to a nitroaromatic compound [Srinivasan et al., 1996]. The goal is to predict the mutagenicity of each compound on Salmonella typhimurium. Nodes in the graphs correspond to atoms (and are associated with a one-hot encoded feature in $\mathbb{R}^{7}$ corresponding to the atom type), and edges correspond to bonds. The average number of nodes in a graph is 17.93, and the average number of edges is 19.79.
@@
The PROTEINS dataset Borgwardt et al. [2005] consists of 1113 graphs, each of which corresponds to a protein. The task is predicting whether or not the protein is an enzyme. Nodes in the graph correspond to amino acids in the protein (and are associated with node features in $\mathbb{R}^{3}$ representing amino acid properties). Edges connect amino acids that are less than some threshold distance from one another in the protein. The average number of nodes is 39.06, and the average number of edges is 72.82.
Peptides-func & Peptides-struct
The peptides-func and peptides-struct datasets Dwivedi et al. [2022] consist of the same 15535 graphs, with different prediction targets. Each graph corresponds to a peptide; nodes correspond to heavy atoms (and are associated with 9 categorical node features representing atom properties) and edges correspond to bonds (and are associated with 3 categorical edge features representing bond properties). For peptides-func, the prediction task is multi-label graph classification (10 classes) of the peptide function. For peptides-struct, the prediction task is graph regression of various peptide properties (11 regression targets). The average number of nodes in a graph is 150.94, and average number of edges is 307.30. We use a train/valid/test split consistent with Dwivedi et al. [2022].
@@
The PPI dataset [Hamilton, 2020] consists of 24 graphs, each of which corresponds to a protein-protein interaction network found in different areas of the body. Each node in the graph corresponds to a protein, with edges connecting proteins that interact with one another. Nodes are associated with features in $\mathbb{R}^{50}$, representing some properties of the protein. Each protein has $121$ binary prediction targets, each of which corresponds to some ontological property that the protein may or may not have. We use a 20/2/2 train/valid/test split consistent with Hamilton et al. [2017].
@@ G.2 Results
For PROTEINS and MUTAG, we perform 10-fold cross validation and report average classification accuracy and standard deviations in Table 3. For PPI, we use a 20/2/2 train/valid/test split consistent with Hamilton et al. [2017], and report average micro-f1 scores in Table 4. For Peptides-func and Peptides-struct, we use a train/valid/test split consistent with Dwivedi et al. [2022] and report average precision and mean average error, respectively, in Table 5. Non-asterisked values correspond to our experimental setup, where the number of parameters (and embedding dimension) is equal across architectures. When parameter numbers are equal, the energy GNN architecture achieves the best performance on MUTAG, PPI, and Peptides-func, and achieves competitive performance on PROTEINS and Peptides-struct. We note that these reported results are under synchronous evaluation (similar to the synthetic experiments, Table 6 demonstrates that under asynchronous execution, we observe a decrease in performance for GCN and GAT, the representative explicitly-defined GNNs).
In Tables 3 and 4, we also include performance reported by other works (marked by an asterisk), which correspond to architectures using a higher number of parameters and larger embedding dimensions. Where layer specifications are not included (for asterisked values), we were unable to determine them from the cited paper. Experiments with larger energy GNN architectures, including multi-layer energy GNN architectures, is left to future work.
@@
Energy GNN (edge-wise) 87.6 $\pm$ 3.4 72.5 $\pm$ 0.3
@@
GCN* [Xu et al., 2019] (5 layer, embedding dimension=64) 85.6 $\pm$ 5.8 76.0 $\pm$ 3.2
IGNN* [Gu et al., 2020] (3 layer, embedding dimension=32) 89.3 $\pm$ 6.7 77.7 $\pm$ 3.4
GIN* [Xu et al., 2019] (5 layer, embedding dimension=64) 89.4 $\pm$ 5.6 76.2 $\pm$ 1.9
Table 3: Graph classification accuracy (%). Results are averaged (and standard deviations are computed) using 10 fold cross validation with 5 random parameter seeds. Asterisked values are obtained from the work cited. Non-asterisked values correspond to architectures with the same number of parameters and 2-dimensional embeddings.
@@
Energy GNN (edge-wise) 76.2
@@
MLP* [Gu et al., 2020] 46.2
GCN* [Gu et al., 2020] 59.2
GraphSAGE* [Veličković et al., 2018] (3 layer, embedding dimensions=[512, 512, 726]) 76.8
GAT* [Veličković et al., 2018] (3 layer, embedding dimension=1024) 97.3
IGNN* [Gu et al., 2020] (5 layer, embedding dimensions=[1024, 512, 512, 256, 121]) 97.6
Table 4: Mean micro-F1 score for node classification on PPI dataset (%). Asterisked values are obtained from the work cited. Non-asterisked values correspond to architectures with the same number of parameters and 2-dimensional embeddings.
DATASET
MODEL Peptides-func (AP) Peptides-struct (MAE)
Energy GNN (edge-wise) 0.348 0.367
Energy GNN + attention 0.381 0.402
IGNN (1 layer) 0.211 0.426
GSD GNN 0.343 0.375
GCN (5 layer) 0.362 0.366
GAT (5 layer) 0.355 0.335
Table 5: Performance on Peptides-func (average precision) and Peptides-struct (mean average error) datasets from the LRGB benchmarks. Values correspond to architectures with the same number of parameters and 2-dimensional embeddings.
DATASET
MODEL MUTAG (%) PROTEINS (%) PPI (micro f1) Peptides-func (AP) Peptides-struct (MAE)
GCN (5 layer) 16.0 $\pm$ 10.2 24.0 $\pm$ 4.9 25.1 $\pm$ 9.5 0.194 $\pm$ 0.085 0.734 $\pm$ 0.041
GAT (5 layer) 28.0 $\pm$ 11.7 40.0 $\pm$ 5.8 29.4 $\pm$ 12.3 0.494 $\pm$ 0.031 0.335 $\pm$ 0.115
Table 6: Decrease in task performance observed from switching from synchronous to asynchronous inference on sub-sample of test data (10 samples) using one trained model instance. Mean and standard deviation are across 5 asynchronous runs. The poor performance of GCN and GAT are consistent with the expected unreliability of explicitly-defined GNNs with asynchronous inference. Decreases in task performance for all implicitly-defined GNNs (IGNN, GSDGNN, and energy GNN variants) is less than $0.1\%$ of synchronous performance (i.e. result from numerical error); these values are omitted from the table for concision.